Opened 7 years ago

Closed 10 months ago

#468 closed flight processing (fixed)

ET12/14, flight day 324/2012, Boset

Reported by: mark1 Owned by: knpa
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Processing: general Keywords:
Cc: Other processors: edfi

Description (last modified by edfi)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/ethiopia/ET12_14-2012_324_Boset

Data arrived from ARSF via SATA disk on 6th December 2012.

Scientific objective: to understand a) how faults initiate, grow and interact during extension, b) the distribution, volume and chronology of hazardous lava flows, and c) the chronology of distribution of faulting and volcanic construction in creating rift valley topography.

Priority: 8

PI: Derek Keir

Sensors:

  • Eagle (03/05/2013)
  • Hawk (03/05/2013)
  • Leica LIDAR (05/02/2013)
  • RCD (05/02/2013)

Change History (33)

comment:1 Changed 7 years ago by mark1

  • Priority changed from alpha 4 high to alpha 4 medium

comment:2 Changed 7 years ago by knpa

Created RCD Tiffs

comment:3 Changed 7 years ago by mark1

GPS data provided by PI copied to basestation directory (15/01/2013)

comment:4 Changed 7 years ago by knpa

Starting nav processing.

comment:5 Changed 7 years ago by knpa

  • Owner set to knpa
  • Status changed from new to assigned

comment:6 Changed 7 years ago by knpa

Navigation complete.

Camera sol file made too.

comment:7 Changed 7 years ago by knpa

The camera cut out during the 3rd line of this flight.

The first 219 images are fine, however images 220-2124 are unusable.

The event file has information for 2886 images.

comment:8 Changed 7 years ago by knpa

  • Milestone set to 2012 data processing completion
  • Other processors set to edfi

comment:9 Changed 7 years ago by edfi

LiDAR Started

comment:10 Changed 7 years ago by knpa

Camera delivery completed.

Ready for delivery check.

comment:11 Changed 7 years ago by edfi

Last edited 7 years ago by edfi (previous) (diff)

comment:12 Changed 7 years ago by edfi

Starting Camera delivery check

comment:13 Changed 7 years ago by edfi

check complete ready to deliver

comment:14 Changed 7 years ago by edfi

Started Hyperspectral processing

comment:15 Changed 7 years ago by knpa

Put camera delivery onto disk 130. Waiting for LiDAR as that is almost finished.

comment:16 Changed 7 years ago by mark1

  • Description modified (diff)
  • Summary changed from ET12/14, flight day 323/2012, Boset to ET12/14, flight day 324/2012, Boset

comment:17 Changed 7 years ago by edfi

LiDAR processing complete awaiting checks

comment:18 Changed 7 years ago by anhi

LiDAR Deliver Check:

  • Dem looks a bit too big to see the LiDAR data in.
  • Some of the elevation offsets in the Readme are repeated: 7 and 8, 14 and 16, 14 and 17.

No other problems

comment:19 Changed 6 years ago by edfi

Changes made

DEM size adjusted, repeated lines removed and line numbers changed

Ready for re-check

comment:20 Changed 6 years ago by anhi

LiDAR Delivery Check:

No problems found.

comment:21 Changed 6 years ago by edfi

Loading LiDAR and Camera onto HDD 324

comment:22 Changed 6 years ago by edfi

LiDAR and Camera Dispatched

comment:23 Changed 6 years ago by edfi

  • Description modified (diff)

comment:24 Changed 6 years ago by edfi

Starting to process eagle using December 2012 calibration

comment:25 Changed 6 years ago by anhi

Hyperspectral Delivery check:

  • Mention dropped scans in eagle and hawk lines 3 and 17
  • Eagle lines 12 and 14 and Hawk line 2 SCTs look a little off

No other problems

comment:26 Changed 6 years ago by edfi

Drop lines mentioned in delivery and re scted line 14

comment:27 Changed 6 years ago by anhi

Hyperspectral Delivery check:

No problems found

comment:28 Changed 6 years ago by edfi

Eagle and Hawk dispatched to PI

comment:29 Changed 6 years ago by edfi

  • Description modified (diff)

comment:30 Changed 5 years ago by dac

Starting to archive dataset.

comment:31 Changed 5 years ago by dac

Submitted to NEODC for archiving.

comment:32 Changed 4 years ago by dac

Processed and raw data uploaded to NEODC via rsync (completed 26/01/2014)

comment:33 Changed 10 months ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed

Archiving

Data available from CEDA. Closing ticket.

Note: See TracTickets for help on using tickets.